# and the Quantum Hole Argument

### v1

@@ VI.1 The Quantum Hole Argument
However, when comparing models that lie on different orbits, i.e. that are not mapped into one another by any diffeomorphism, there is an additional ambiguity that comes from the choice of QRF. In particular, this is true even when using coincidences of physical field values to characterise the spacetime points, in the spirit of Einstein’s “point-coincidence” argument [66, 77]. For even when using such a gauge-invariant characterisation of spacetime points, there is still the question of which fields to use in order to define the coincidences. This corresponds to the freedom of choosing either the $\chi-$ or the $\tilde{\chi}-$ fields as QRFs in Sec. V above.
@@
Figure 12: Action of step (i) of a QRF transformation, i.e. a quantum diffeomorphism, on a superposition of models $(\mathcal{M}_{1},g_{ab}^{(1)},\chi_{(A)}^{(1)})$ and $(\mathcal{M}_{2},g_{ab}^{(2)},\chi_{(A)}^{(2)})$ with the comparison map $\mathrm{C}_{\chi}$ established through the $\chi$-fields.
In particular, with respect to $\mathrm{C}_{\chi}$, the point $p\in\mathcal{M}_{1}$ is identified with the point $q=\mathrm{C}_{\chi}(p)\in\mathcal{M}_{2}$. We say that the pair $(p,q)$ is localised with respect to the $\chi$-fields. The quantum diffeomorphism consists of $d^{(1)}$ and $d^{(2)}$ in the two branches of the superposition and maps to the models $(\mathcal{M}_{1},{g^{\prime}}_{ab}^{(1)},{\chi^{\prime}}_{(A)}^{(1)})$ and $(\mathcal{M}_{2},{g^{\prime}}_{ab}^{(2)},{\chi^{\prime}}_{(A)}^{(2)})$, respectively. It changes the comparison map $\mathrm{C}_{\chi}$ to $\mathrm{C}_{\chi}^{\prime}$ in such a way that the pair $(p,q)$ remains localised relative to the new comparison map $\mathrm{C}_{\chi}^{\prime}$, specifically, $d^{(2)}(q)=\mathrm{C}_{\chi}^{\prime}(d^{(1)}(p))$. In this sense, the quantum diffeomorphism leaves the identification of points intact. Note that step (ii) of the QRF transformation, not depicted in this figure, consists of a change of section, i.e. a change to the corresponding quantum reference fields $\tilde{\chi}$. This step changes the comparison map to $\mathrm{C}_{\tilde{\chi}}$, which will, in general, no longer identify $d^{(1)}(p)$ and $d^{(2)}(q)$: $d^{(2)}(q)\neq\mathrm{C}_{\tilde{\chi}}(d^{(1)}(p))$.
@@
To illustrate this observation, let us consider a concrete example. The idea here is to take four scalars $R_{(A)},\,\,A=1,\dots,4$, formed by certain real scalar functions of the Riemann tensor $R^{a}_{\;\;bcd}$ (cf. [67, 68] and [32, Sec. 5.1.b]). [^25] Komar [68] finds these real scalars through an eigenvalue problem: $(R_{{a}{b}{c}{d}}-\lambda(g_{{a}{c}}g_{{b}{d}}+g_{{a}{d}}g_{{b}{c}}))V^{{c}{d}}=0,$ where $V^{{c}{d}}$ is an anti-symmetric tensor. The requirement ensures that solutions $\lambda$, whose existence we assume, are independent real scalar functions. While Komar takes these scalars to be preferred, since “they are the only nontrivial scalars which are of least possible order in derivatives of the metric” [68, p. 1183], we do not need to take the simplest choices. Nonetheless, to keep our notation tractable, we will use sets that have the next-to-lowest order of derivatives (namely, one in which we take the D’Alembertian of one curvature scalar). The reason we take $\square R$, and not $R$, and similarly with $\mathrm{Weyl}^{2}$, is so that we do not have to keep track of the following functional dependence relation: $\mathrm{Weyl}^{2}=\mathrm{Riem}^{2}-2\mathrm{Ric}^{2}+\frac{1}{3}R^{2}$ Suppose that we choose
@@ VI.2 Localisation of Events, Interference, and Recombination
At first sight, this may raise worries when considering an interaction that is defined to take place at a localised event, since that event can be mapped to a delocalised one through a change of QRF. Thus, Adlam, Linnemann, and Read have argued: “it would seem that using different identities across branches would produce different predictions, since we would end up with different combinations of fields interacting” [62, p. 13, cf. also pp. 33-36]. However, while a QRF transformation may change the localisation of an event at which the interaction is taking place, it does not change the locality of the interaction itself. The latter is a notion defined separately in each branch. It simply refers to whether the fields interact at a single point within a given spacetime, not whether it is the same point across the different branches. Therefore, the interaction of the fields remains the same in each branch, with no empirical consequences upon changes of QRFs. While this only concerns the interaction within a given spacetime, Adlam et al. are particularly concerned with phenomena occurring across branches, such as interference or recombination:
“Consider for illustration the BMV experiment [our Refs. [74, 75]] in which two masses in superposition states are taken to get entangled with each other through gravitational interaction (and gravitational interaction alone): the relevant literature tacitly assumes that the location of the masses are all relative to one joint lab frame—no matter whether the experiment is modelled through a Newtonian potential, or (low-energy) metric fields, as done by Christodoulou and Rovelli (2019) [our Ref. [64]]. But if we perform a quantum diffeomorphism which shifts the point at which the recombination occurs within one of the branches, then the phase change will be different and the interference effects will change.” [62, p. 13]
@@ VI.3 Relational Observables
To get a more encompassing understanding of the effects of changing the QRF and the identification of points, let us further consider its implications for observables. Let us denote a partial observable [78] on a classical spacetime by a function $\mathcal{O}[g_{ab},\Psi_{\mathrm{matter}}]:\mathcal{M}\to\mathcal{S}$ that depends on the metric and matter fields as well as arbitrary derivatives thereof. For simplicity, we restrict ourselves in the following discussion to scalar functions for which $\mathcal{S}=\mathbb{R}$. On a superposition of spacetimes, a partial observable is then given by a collection of classical partial observables, $\mathbf{O}:=\{\mathcal{O}^{(i)}\equiv\mathcal{O}[g_{ab}^{(i)},\Psi_{\mathrm{matter}}^{(i)}]\}_{i}$ where $i$ denotes the respective branch of the superposition. Given a partial observable and a choice of reference frame $\chi=\chi_{\ast}$ with corresponding section $\sigma$, one can define the relational or dressed observable $\mathbf{O}_{|\chi=\chi_{\ast}}:=\{\mathcal{O}^{(i)}_{|\chi^{(i)}=\chi_{\ast}}\}_{i}:=\{\mathcal{O}^{(i)}\circ\left(\chi^{(i)}\right)^{-1}|_{\chi^{(i)}=\chi_{\ast}}\}_{i}$. [^27] Note that, when using idealised reference fields (in the sense of the first of the three options listed at the start of Sec. V), these relational observables can still capture any possible configuration (state) of the metric relative to these reference fields. When using dynamically coupled reference fields (in the sense of either the second or the third of the three options above), on the other hand, fixing the values of the reference fields also leads to a unique gauge fixing for the spacetime geometry. This distinction will be discussed further in [79]. This is a set of functions from $\mathbb{R}^{4}$ (understood as the range of the four $\chi$-fields, taken together), back to the respective spacetime manifolds $\mathcal{M}_{i}$, and then to the value-space $\mathcal{S}$, where the $i$-th function gives the value of the partial observable $\mathcal{O}^{(i)}$ when the reference field $\chi^{(i)}$ takes the value $\chi_{\ast}$. These relational or dressed observables have not only been studied in the context of QRFs [7, 14, 70] but also play an important role in certain approaches to quantum gravity [80, 78, 81, 82, 71]. An example of a partial observable would be the curvature scalar $\mathbf{R}$, which takes a point in $\mathcal{M}_{i}$ and returns the value of the curvature scalar for the metric $g_{ab}^{(i)}$ on each manifold in the superposition. When choosing a reference frame $\chi=\chi_{\ast}$, we get the corresponding relational observable $\mathbf{R}_{|\chi=\chi_{\ast}}$ which is a set of maps from the space $\mathbb{R}^{4}$ of values for the $\chi$-fields to $\mathcal{S}=\mathbb{R}$. Here is a concrete example: given a value for the scalar fields, $x\in\mathbb{R}^{4}$, it returns the curvature scalar $R^{(i)}(p_{i})=R^{(i)}((\chi_{\ast}^{(i)})^{-1}(x))$ for $p_{i}=((\chi_{\ast}^{(i)})^{-1})(x)$ on each manifold in the superposition.
@@ VI.4 Relation to Indefinite Causal Order
Let us briefly introduce the problem. One prime example of a process with indefinite causal order is the quantum switch [42, 43], in which two agents apply operations on a target system in a superposition of opposite orders. The application of each operation marks an event, often represented as the target system passing through the laboratory of the respective agent. There are different implementations of the quantum switch. In the optical quantum switch [44], the indefinite causal order arises through a superposition of paths of the target system on a definite Minkowski spacetime. In the gravitational quantum switch [83], it is realised through a gravitational field in superposition. When sketching both processes in a spacetime diagram, the optical switch is depicted as a “four-point” switch in which the target system passes through the laboratories at four different spacetime points while the gravitational implementation can be constructed as a “two-point” switch [46].
There is an ongoing debate on whether a proper realisation of indefinite causal order requires a superposition of spacetimes or whether an implementation in Minkowski spacetime, such as the optical switch, is sufficient [44, 45, 46, 84, 47]. A crucial question in this debate is how many events there are in the different implementations of the quantum switch [46, 47, 85, 86, 84]. The answer to this question of course depends largely on how one defines a quantum event. In [19], some of us proposed a notion of event based on worldline coincidences which encompasses situations where both the trajectories and the spacetimes can be in superposition. Given a superposition of models $\varphi_{i}$, $i=1,2$, each encoding the configuration of the spacetime metric as well as several timelike worldlines, we defined an event by the pair of points $p\in\mathcal{M}_{1}$ and $q\in\mathcal{M}_{2}$ at which the worldlines cross in the respective manifolds. To relate the different implementations of the quantum switch, we then applied a quantum-controlled transformation that maps a four-point to a two-point switch, thereby localising the events.
@@ VII Conclusion
Both of these steps are crucial to the conceptual understanding of some of the key features of QRF transformations. Only if we change the configurations through a quantum-controlled symmetry transformation and change our way of identifying configurations or locations across the branches of a superposition, can we ensure that the state of the reference frame factorises out — i.e. is “the same” across all branches — and alter whether a system is in a superposition — that is, in “different” configurations across the branches. Once one understands that the way we compare objects across a superposition is not pre-defined but depends crucially on the choice of QRF, the frame-dependence of both superposition and entanglement finds a natural explanation. While the relativity of sameness and difference can already be understood at the classical level when comparing different possible configurations, it has more physical implications when one considers quantum superpositions. Similarly, in the case of ideal QRFs, the relativity of subsystems [2, 13, 10, 23] can, maybe surprisingly, already be understood classically through a reshuffling of the degrees of freedom (cf. [23]), but becomes particularly relevant when studying entanglement in quantum theory. Note, however, that it is still an open question whether subsystem relativity for non-ideal frames, that is, frames that do not carry a representation in $L^{2}(G)$ [10], can be captured by this explanation.
@@
While our framework is mainly based on the understanding of QRFs as in [2, 5, 8, 11, 17, 18], it can be connected to many other approaches towards making physics relational. Firstly, visualising QRF transformations in the space of models makes clear that they have a classical analogue — the model- or configuration-dependent symmetry transformations. When considering field theories, these become field-dependent maps, such as the changes between dynamical reference frames in [87, 70, 88]. These, in turn, are directly related to edge modes, additional degrees of freedom that arise when considering a bounded region of spacetime in gauge theory and gravity (e.g. [89, 90, 91]). As shown in [92, 93, 94, 21], edge modes can be understood as dynamical reference frames and used to construct relational observables. On the quantum level, they can be seen as QRFs and define a relational evolution of the quantum state for generally covariant theories. When constructing relational observables, both edge modes and dynamical reference frames act as dressing functions that turn the bare observables into gauge-invariant quantities relative to the reference fields [93]. This brings us back to the space of models, where the group elements $g_{\sigma}$ that map the models to a chosen section $\sigma$ can equally be seen as dressings for the models [26].
Secondly, while the connection of the present framework to QRF approaches based on observable algebras [1, 22] remains an open question, it is possible to connect it to some aspects of the perspective-neutral approach to QRFs [3, 7, 14, 23]. It has been shown [14] that the latter is equivalent to the approach of [2, 8] for ideal QRFs, so our framework carries over directly in this context. Moreover, as discussed in Sec. IV, the present work encompasses non-ideal QRFs with non-trivial stabiliser groups, as developed in [14]. What is still to be worked out is the formal treatment of non-ideal QRFs using coherent states (cf. the end of Sec. IV).

### v2

@@ VI.1 The Quantum Hole Argument
However, when comparing models that lie on different orbits, i.e. that are not mapped into one another by any diffeomorphism, there is an additional ambiguity that comes from the choice of QRF. In particular, this is true even when using coincidences of physical field values to characterise the spacetime points, in the spirit of Einstein’s “point-coincidence” argument [69, 80]. For even when using such a gauge-invariant characterisation of spacetime points, there is still the question of which fields to use in order to define the coincidences. This corresponds to the freedom of choosing either the $\chi-$ or the $\tilde{\chi}-$ fields as QRFs in subsection Sec. V above.
@@
Figure 12: The first step of a QRF transformation: the application of a quantum diffeomorphism. The quantum diffeomorphism acts on a superposition of models $(\mathcal{M}_{1},g_{ab}^{(1)},\chi_{(A)}^{(1)})$ and $(\mathcal{M}_{2},g_{ab}^{(2)},\chi_{(A)}^{(2)})$ with the comparison map $\mathrm{C}_{\chi}$ established through the $\chi$-fields.
In particular, with respect to $\mathrm{C}_{\chi}$, the point $p\in\mathcal{M}_{1}$ is identified with the point $q=\mathrm{C}_{\chi}(p)\in\mathcal{M}_{2}$. We say that the pair $(p,q)$ is localised with respect to the $\chi$-fields. The quantum diffeomorphism consists of $d^{(1)}$ and $d^{(2)}$ in the two branches of the superposition and maps to the models $(\mathcal{M}_{1},{g^{\prime}}_{ab}^{(1)},{\chi^{\prime}}_{(A)}^{(1)})$ and $(\mathcal{M}_{2},{g^{\prime}}_{ab}^{(2)},{\chi^{\prime}}_{(A)}^{(2)})$, respectively. It changes the comparison map $\mathrm{C}_{\chi}$ to $\mathrm{C}_{\chi}^{\prime}$ in such a way that the pair $(p,q)$ remains localised relative to the new comparison map $\mathrm{C}_{\chi}^{\prime}$, specifically, $d^{(2)}(q)=\mathrm{C}_{\chi}^{\prime}(d^{(1)}(p))$. In this sense, the quantum diffeomorphism leaves the identification of points intact. Note that a full QRF transformation involves of a second step, not depicted in this figure, namely a change of section, i.e. a change to the corresponding quantum reference fields $\tilde{\chi}$. This step changes the comparison map to $\mathrm{C}_{\tilde{\chi}}$, which will, in general, no longer identify $d^{(1)}(p)$ and $d^{(2)}(q)$: $d^{(2)}(q)\neq\mathrm{C}_{\tilde{\chi}}(d^{(1)}(p))$.
@@
To illustrate this observation, let us consider a concrete example. The idea here is to take four scalars $R_{(A)},\,\,A=1,\dots,4$, formed by certain real scalar functions of the Riemann tensor $R^{a}_{\;\;bcd}$ (cf. [70, 71] and [32, Sec. 5.1.b]). [^27] Komar [71] finds these real scalars through an eigenvalue problem: $(R_{{a}{b}{c}{d}}-\lambda(g_{{a}{c}}g_{{b}{d}}+g_{{a}{d}}g_{{b}{c}}))V^{{c}{d}}=0,$ where $V^{{c}{d}}$ is an anti-symmetric tensor. The requirement ensures that solutions $\lambda$, whose existence we assume, are independent real scalar functions. While Komar takes these scalars to be preferred, since “they are the only nontrivial scalars which are of least possible order in derivatives of the metric” [71, p. 1183], we do not need to take the simplest choices. Nonetheless, to keep our notation tractable, we will use sets that have the next-to-lowest order of derivatives (namely, one in which we take the D’Alembertian of one curvature scalar). The reason we take $\square R$, and not $R$, and similarly with $\mathrm{Weyl}^{2}$, is so that we do not have to keep track of the following functional dependence relation: $\mathrm{Weyl}^{2}=\mathrm{Riem}^{2}-2\mathrm{Ric}^{2}+\frac{1}{3}R^{2}$ Suppose that we choose
@@ VI.2 Localisation of Events, Interference, and Recombination
At first sight, this may raise worries when considering an interaction that is defined to take place at a localised event, since that event can be mapped to a delocalised one through a change of QRF. Thus, Adlam, Linnemann, and Read have argued: “it would seem that using different identities across branches would produce different predictions, since we would end up with different combinations of fields interacting” [66, p. 13, cf. also pp. 33-36]. However, while a QRF transformation may change the localisation of an event at which the interaction is taking place, it does not change the locality of the interaction itself. The latter is a notion defined separately in each branch. It simply refers to whether the fields interact at a single point within a given spacetime, not whether it is the same point across the different branches. Therefore, the interaction of the fields remains the same in each branch, with no empirical consequences upon changes of QRFs. While this only concerns the interaction within a given spacetime, Adlam et al. are particularly concerned with phenomena occurring across branches, such as interference or recombination:
“Consider for illustration the BMV experiment [our Refs. [77, 78]] in which two masses in superposition states are taken to get entangled with each other through gravitational interaction (and gravitational interaction alone): the relevant literature tacitly assumes that the location of the masses are all relative to one joint lab frame—no matter whether the experiment is modelled through a Newtonian potential, or (low-energy) metric fields, as done by Christodoulou and Rovelli (2019) [our Ref. [68]]. But if we perform a quantum diffeomorphism which shifts the point at which the recombination occurs within one of the branches, then the phase change will be different and the interference effects will change.” [66, p. 13]
@@ VI.3 Relational Observables
To get a more encompassing understanding of the effects of changing the QRF and the identification of points, let us further consider its implications for observables. Let us denote a partial observable [81] on a classical spacetime by a function $\mathcal{O}[g_{ab},\Psi_{\mathrm{matter}}]:\mathcal{M}\to\mathcal{S}$ that depends on the metric and matter fields as well as arbitrary derivatives thereof. For simplicity, we restrict ourselves in the following discussion to scalar functions for which $\mathcal{S}=\mathbb{R}$. On a superposition of spacetimes, a partial observable is then given by a collection of classical partial observables, $\mathbf{O}:=\{\mathcal{O}^{(i)}\equiv\mathcal{O}[g_{ab}^{(i)},\Psi_{\mathrm{matter}}^{(i)}]\}_{i}$ where $i$ denotes the respective branch of the superposition. Given a partial observable and a choice of reference frame $\chi=\chi_{\ast}$ with corresponding section $\sigma$, one can define the relational or dressed observable $\mathbf{O}_{|\chi=\chi_{\ast}}:=\{\mathcal{O}^{(i)}_{|\chi^{(i)}=\chi_{\ast}}\}_{i}:=\{\mathcal{O}^{(i)}\circ\left(\chi^{(i)}\right)^{-1}|_{\chi^{(i)}=\chi_{\ast}}\}_{i}$. [^29] Note that, when using idealised reference fields (in the sense of the first of the three options listed at the start of subsection Sec. V), these relational observables can still capture any possible configuration (state) of the metric relative to these reference fields. When using dynamically coupled reference fields (in the sense of either the second or the third of the three options above), on the other hand, fixing the values of the reference fields also leads to a unique gauge fixing for the spacetime geometry. This distinction is discussed further in [82]. This is a set of functions from $\mathbb{R}^{4}$ (understood as the range of the four $\chi$-fields, taken together), back to the respective spacetime manifolds $\mathcal{M}_{i}$, and then to the value-space $\mathcal{S}$, where the $i$-th function gives the value of the partial observable $\mathcal{O}^{(i)}$ when the reference field $\chi^{(i)}$ takes the value $\chi_{\ast}$. These relational or dressed observables have not only been studied in the context of QRFs [7, 14, 73] but also play an important role in certain approaches to quantum gravity [83, 81, 84, 85, 74]. An example of a partial observable would be the curvature scalar $\mathbf{R}$, which takes a point in $\mathcal{M}_{i}$ and returns the value of the curvature scalar for the metric $g_{ab}^{(i)}$ on each manifold in the superposition. When choosing a reference frame $\chi=\chi_{\ast}$, we get the corresponding relational observable $\mathbf{R}_{|\chi=\chi_{\ast}}$ which is a set of maps from the space $\mathbb{R}^{4}$ of values for the $\chi$-fields to $\mathcal{S}=\mathbb{R}$. Here is a concrete example: given a value for the scalar fields, $x\in\mathbb{R}^{4}$, it returns the curvature scalar $R^{(i)}(p_{i})=R^{(i)}((\chi_{\ast}^{(i)})^{-1}(x))$ for $p_{i}=((\chi_{\ast}^{(i)})^{-1})(x)$ on each manifold in the superposition.
@@ VI.4 Relation to Indefinite Causal Order
Let us briefly introduce the problem. One prime example of a process with indefinite causal order is the quantum switch [42, 43], in which two agents apply operations on a target system in a superposition of opposite orders. The application of each operation marks an event, often represented as the target system passing through the laboratory of the respective agent. There are different implementations of the quantum switch. In the optical quantum switch [44, 86], the indefinite causal order arises through a superposition of paths of the target system on a definite Minkowski spacetime. In the gravitational quantum switch [87], it is realised through a gravitational field in superposition. When sketching both processes in a spacetime diagram, the optical switch is depicted as a “four-point” switch in which the target system passes through the laboratories at four different spacetime points while the gravitational implementation can be constructed as a “two-point” switch [46].
There is an ongoing debate on whether a proper realisation of indefinite causal order requires a superposition of spacetimes or whether an implementation in Minkowski spacetime, such as the optical switch, is sufficient [44, 45, 46, 88, 47, 89]. A crucial question in this debate is how many events there are in the different implementations of the quantum switch [46, 47, 90, 91, 88]. The answer to this question of course depends largely on how one defines a quantum event. In [19], some of us proposed a notion of event based on worldline coincidences which encompasses situations where both the trajectories and the spacetimes can be in superposition. Given a superposition of models $\varphi_{i}$, $i=1,2$, each encoding the configuration of the spacetime metric as well as several timelike worldlines, we defined an event by the pair of points $p\in\mathcal{M}_{1}$ and $q\in\mathcal{M}_{2}$ at which the worldlines cross in the respective manifolds. To relate the different implementations of the quantum switch, we then applied a quantum-controlled transformation that maps a four-point to a two-point switch, thereby localising the events.
@@ VII Conclusion
Both of these steps are crucial to the conceptual understanding of some of the key features of QRF transformations. Only if we change the configurations through a quantum-controlled symmetry transformation and change our way of identifying configurations or locations across the branches of a superposition, can we ensure that the state of the reference frame factorises out — i.e. is “the same” across all branches — and alter whether a system is in a superposition — that is, in “different” configurations across the branches. Once one understands that the way we compare objects across a superposition is not pre-defined but depends crucially on the choice of QRF, the frame-dependence of both superposition and entanglement finds a natural explanation. While the relativity of sameness and difference can already be understood at the classical level when comparing different possible configurations, it has more physical implications when one considers quantum superpositions. Similarly, in the case of ideal QRFs, the relativity of subsystems [2, 13, 10, 14, 23] can, maybe surprisingly, already be understood classically through a reshuffling of the degrees of freedom (cf. [23]), but becomes particularly relevant when studying entanglement in quantum theory. Note, however, that it is still an open question whether subsystem relativity for non-ideal frames, that is, frames that do not carry a representation in $L^{2}(G)$ [10], can be captured by this explanation.
@@
One of the most pressing problems is, of course, the rigorous formalisation of both the measures over the diffeomorphism group and the space of metrics as well as the construction of a well-defined Hilbert space for the spacetime metric. While some important advances have been made in this regard, e.g. in the context of lattice and algebraic quantum field theory or spin foams [92, 93, 94], these problems have remained unsolved for several decades. We believe that this should not hinder our efforts to understand some of the novel features of non-classical spacetimes with the tools that we do have on hand. It is true, of course, that we cannot exclude the possibility that a rigorous solution to the aforementioned problems might reveal unforeseen effects and it is therefore important to find a balance between addressing the known challenges and making progress with current techniques. Moreover, advancements in both areas can complement each other and thus guide the development of a more complete theory in the future.
Moreover, while our framework is mainly based on the understanding of QRFs as in [2, 5, 8, 11, 17, 18], it can be connected to many other approaches towards making physics relational. Firstly, visualising QRF transformations in the space of models makes clear that they have a classical analogue — the model- or configuration-dependent symmetry transformations. When considering field theories, these become field-dependent maps, such as the changes between dynamical reference frames in [95, 73, 96]. These, in turn, are directly related to edge modes, additional degrees of freedom that arise when considering a bounded region of spacetime in gauge theory and gravity (e.g. [97, 98, 99]). As shown in [100, 101, 102, 21], edge modes can be understood as dynamical reference frames and used to construct relational observables. On the quantum level, they can be seen as QRFs and define a relational evolution of the quantum state for generally covariant theories. When constructing relational observables, both edge modes and dynamical reference frames act as dressing functions that turn the bare observables into gauge-invariant quantities relative to the reference fields [101]. This brings us back to the space of models, where the group elements $g_{\sigma}$ that map the models to a chosen section $\sigma$ can equally be seen as dressings for the models [26].
Secondly, while the connection of the present framework to QRF approaches based on observable algebras [1, 22] as well as the framework introduced in [13] remains an open question, it is possible to connect it to some aspects of the perspective-neutral approach to QRFs [3, 7, 14, 23]. It has been shown [14] that the latter is equivalent to the approach of [2, 8] for ideal QRFs, so our framework carries over directly in this context. Moreover, as discussed in Sec. IV, the present work encompasses non-ideal QRFs with non-trivial stabiliser groups, as developed in [14]. What is still to be worked out is the formal treatment of non-ideal QRFs using coherent states (cf. the end of Sec. IV). Finally, recent advances in the study of gravitational observable algebras [103, 104, 105, 106, 107, 108, 109, 110, 111] have raised important questions about their connection to quantum reference frames [112, 113, 49, 114, 50].
@@
Note added: At around the same time of this article’s first appearance as a pre-print, a paper [117] came out with similar conclusions about the relative localisation of events in the context of linearised general relativity and Jackiw-Teitelboim gravity. The overlap between these approaches clearly deserves further study.
Author contributions
VK, ACdlH, LA, CC, HG, JB, and ČB contributed to all aspects of the research, with leading and equal input of VK, ACdlH, and LA.
Competing interests
The authors declare no competing interests.
Data Availability
Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.